OpenStudy (anonymous):

How that if u increase R by 1% u get 0.01r

OpenStudy (anonymous):

1.01

OpenStudy (anonymous):

yes, you will get an additional of 0.01r plus the r is equal to 1.01r

OpenStudy (anonymous1):

Actually you get 1.01r (or 101% of r), like kumarsajan said. r + 0.01r = 1.01r.

OpenStudy (anonymous):

\[\LARGE R(1+1\%)\] \[\LARGE R+R\cdot \frac{1}{100}\] \[\LARGE R+\frac{R}{100}\] \[\LARGE \frac{100R}{100 }+\frac{R}{100}\] \[\LARGE \frac{101R}{100 }=1.01R\]

OpenStudy (anonymous1):

Noticing that 1 = 100% = 100/100 and 1% = 0.01 = 1/100.

OpenStudy (anonymous):

ok